# Image Segmentation with Large Language Models: A Survey with Perspectives for Intelligent Transportation Systems

### v1

@@ VII-A Autonomous Driving and Scene Understanding
The most prominent application of VLSeg is in enhancing the perception systems of autonomous vehicles. Accurate, real-time scene understanding is the bedrock of safe navigation. LLM-augmented systems allow for a level of semantic richness that was previously unattainable. For instance, a system can be prompted to ”segment the road surface, but exclude any wet patches or oil slicks,” which is a complex instruction that goes beyond simple class labels. This capability is critical for path planning, especially under adverse conditions. Furthermore, models like DriveLM [23] and InsightGPT [24] demonstrate how LLMs can be used not just for segmentation but for integrated reasoning, allowing a vehicle to connect its perception to its planning module (e.g., ”A pedestrian is near the crosswalk, I should yield”). This extends to dynamic obstacle detection, where VLSeg can be used to track vulnerable road users (e.g., cyclists, pedestrians) with high precision, even when they are partially occluded, by using contextual prompts [1, 2].
@@ VII-B Traffic Management and Monitoring
Beyond individual vehicles, VLSeg offers powerful tools for city-scale traffic management. Smart city initiatives rely on networks of cameras to monitor traffic flow. LLM-augmented segmentation can significantly improve the analysis of this data. For example, traffic operators can issue natural language queries to the system, such as ”show me all the trucks that are blocking the intersection at 5th and Main” or ”count the number of vehicles turning left at this junction.” This allows for much more flexible and responsive traffic monitoring than traditional systems that can only detect a predefined set of vehicle classes [25]. These techniques can be used for anomaly detection, such as identifying a stalled vehicle, the formation of an unusual queue, or performing fine-grained temporal localization of crash events in surveillance footage, enabling faster incident response [79, 92]. Frameworks like TrafficGPT [26] are being explored to create conversational interfaces for traffic management, making system operation more intuitive.
@@ VII-C Infrastructure Inspection and Maintenance
A crucial but often overlooked aspect of ITS is the maintenance of the transportation infrastructure itself. VLSeg, guided by LLMs, presents a highly efficient solution for automating the inspection of roads, bridges, and signage. Municipal vehicles equipped with cameras can continuously scan their surroundings. An LLM-based system can then be prompted to find and segment specific types of infrastructure defects. For example, a query like ”segment all potholes deeper than two inches” or ”highlight any traffic signs with visible graffiti or fading” can automate a process that is currently manual, labor-intensive, and slow. This extends to critical pedestrian infrastructure, where robust and precise sidewalk detection is essential for both accessibility and curb management [93]. This is an area of growing research, with datasets emerging that focus specifically on road surface defects and infrastructure anomalies [94, 95]. Automating this process allows for proactive maintenance, improving safety and reducing long-term repair costs.
@@ VII-D Enhanced Urban Mobility and User Experience
VLSeg can also improve the experience of users within the transportation system, including public transit riders and pedestrians. For public transportation, segmentation can be used to monitor passenger flow at bus stops or train stations, or to ensure that dedicated bus lanes are clear of obstructions. For pedestrians, especially those with visual impairments, mobile applications could use VLSeg to provide real-time auditory feedback about the environment, such as ”there is a crosswalk 10 feet ahead to your left” or ”warning: an e-scooter is approaching on the sidewalk.” This creates a more accessible and safer urban environment for everyone. Systems could also enhance navigation services by providing more descriptive guidance, for instance, ”turn right after the large red building,” using segmentation to identify the landmark described [25, 26].
@@ VIII End-to-End Systems and Integrated Reasoning
Sensor DataVLSeg ModuleLLM ReasoningCoreHuman CommandPlanning &ControlVehicleActuatorsHuman-MachineInterfaceSegmented SceneGoalAction CommandExplanationPhysical World
@@ IX-B Remote Assistance and Teleoperation
When an autonomous vehicle encounters a situation it cannot resolve—for example, complex hand gestures from a traffic police officer or an unusual construction zone—it can request help from a remote human operator. VLSeg is crucial for creating an efficient interface for this tele-assistance. The vehicle can stream its sensor data to the operator, who sees a 3D reconstruction of the scene. The operator can then interact with this scene using language and gestures. For example, they could draw a path on the screen and command, ”It is safe to follow this path,” or circle a group of people and ask, ”What are these people doing?” The VLSeg system on the vehicle interprets these multimodal prompts from the operator to navigate safely. Models like Talk2BEV [102] are developing methods to ground these natural language commands from a remote user directly into the Bird’s-Eye-View (BEV) representation used for vehicle planning.
@@ IX-C Driver-AI Collaboration and Shared Autonomy
In vehicles with advanced driver-assistance systems (ADAS) or partial autonomy (SAE Levels 2-3), the driver and the AI are co-pilots. VLSeg can create a much more intuitive and less intrusive collaboration between them. Instead of relying on beeps and cryptic dashboard icons, the vehicle can communicate using language and augmented reality. For example, the system could overlay a segmentation mask on the windshield’s heads-up display and say, ”I see a potential hazard on the right,” highlighting a pedestrian partially obscured by a parked car. Conversely, the driver could interact with the AI using language and gestures. A driver could point to an empty parking space and say, ”Park the car there.” The AI would use VLSeg to precisely segment the indicated space and then execute the parking maneuver. This shared autonomy, explored in frameworks like Voyager [103], aims to make the driving experience safer and more seamless by leveraging the complementary strengths of human intuition and AI perception.
@@ X-B Datasets for Vision-Language and Interactive Segmentation
  • LISA (Language-guided Instance Segmentation) [76]: While not specific to ITS, LISA is a large-scale dataset for reasoning segmentation, where the model must segment objects based on complex queries that require reasoning (e.g., ”segment the car that is farthest away”). This is crucial for developing more intelligent VLSeg systems.
@@ XI Challenges and Future Research Directions
XI-A Challenge 1: Computational Efficiency for Real-Time Systems
@@
XI-B Challenge 2: Data Availability and Open-World Generalization
@@
XI-C Challenge 3: Safety, Reliability, and Explainability
Perhaps the most significant barrier to adoption, and a core perspective of this survey, is the challenge of ensuring safety, reliability, and explainability. Models must be robust against adversarial attacks and perform reliably under adverse conditions. Critically, as foreshadowed in our discussion of end-to-end systems in Section VIII, their decisions must be interpretable, especially in case of failure. Without a clear ”why” behind an action, true safety is unattainable.
@@ XI-C Challenge 3: Safety, Reliability, and Explainability
Future Outlook: Research is moving towards formal verification methods and runtime monitoring to provide safety guarantees [119]. Adversarial training and certified defenses are key areas of research to improve robustness against malicious inputs [120]. Frameworks like Multi-Shield [34] are exploring multimodal defenses. To address safety in high-risk scenarios and provide end-to-end guarantees, recent models like SafeSeg [121] and VLM-AD [122] integrate predictive reasoning with segmentation to directly inform safer driving decisions. A major future direction, and the one most enabled by LLMs, is explainable AI (XAI). The goal is to create systems that not only perform segmentation but also provide causal reasons for their decisions (e.g., ”The object is segmented as a pedestrian because it has human-like shape and motion”), as explored in [123, 124]. Fail-safe mechanisms, including multi-modal redundancy with LiDAR and radar, will be standard [125, 58].
XI-D Challenge 4: System Integration and Standardization
@@
XI-E Challenge 5: Ethical Considerations and Algorithmic Bias
@@
Beyond technical hurdles, the deployment of LLM-augmented segmentation in ITS raises significant ethical questions. The data used to train these models can contain inherent biases, which can lead to inequitable and unsafe outcomes. For example, if a dataset is predominantly collected in one geographic region (e.g., North America), the models may perform worse at recognizing pedestrian behaviors or road signs in other parts of the world. There is a documented ”long-tail” problem where models perform poorly on underrepresented groups [127], which in a driving context, could mean a higher risk for certain demographics of pedestrians. Furthermore, the decision-making process of these large models is often opaque (the ”black box” problem), making it difficult to audit or explain failures, which is a major concern for accountability in the event of an accident [123].
@@
Future Outlook: The ITS research community must prioritize the development of ”fair” and ”transparent” AI. This involves creating more geographically and demographically balanced datasets and developing techniques for bias detection and mitigation [128]. Explainable AI (XAI) is a critical research frontier, aiming to create models that can articulate the reasoning behind their predictions (e.g., ”I am slowing down because I have segmented a child running towards the street”). Future regulatory frameworks will likely mandate auditable AI systems in autonomous vehicles, requiring a shift away from purely performance-driven metrics towards a more holistic evaluation that includes fairness, transparency, and ethical alignment.
XI-F Long-Term Speculative Directions
Looking further ahead, several emerging technologies could reshape VLSeg in ITS. Neuromorphic computing, with its brain-inspired, event-based processing, promises unparalleled energy efficiency for real-time tasks. Hybrid neural-symbolic AI could integrate common-sense reasoning into perception models, allowing them to understand context in a more human-like way. While highly speculative, quantum computing could eventually offer breakthroughs in solving complex optimization problems inherent in large-scale model training and scene understanding. These directions, while not immediately deployable, represent exciting long-term frontiers for the field.

### v2

@@ VII Applications to Intelligent Transportation Systems
a) Autonomous Driving Input Image: Street view with a cyclist Output Mask: Cyclist is segmented Prompt: “highlight the cyclist on the right” b) Traffic Management Input Image: Overhead view of an intersection Output Mask: Trucks are segmented Prompt: “show trucks blocking the intersection” c) Infrastructure Inspection Input Image: Close-up of a road surface Output Mask: Potholes are segmented Prompt: “highlight potholes deeper than two inches” d) Urban Mobility / Accessibility Input Image: Pedestrian view of a street corner Output Mask: Crosswalk is segmented Prompt: “identify the crosswalk ahead”
Figure 7: Application scenarios of Vision-Language Segmentation (VLSeg) in Intelligent Transportation Systems (ITS). Each panel illustrates how a natural language prompt can be used to segment specific objects of interest in a real-world scene, highlighting the practical utility of VLSeg for autonomous driving, traffic management, infrastructure inspection, and urban mobility.
@@ VII-A Autonomous Driving and Scene Understanding
The most prominent application of VLSeg is in enhancing the perception systems of autonomous vehicles. Accurate, real-time scene understanding is the bedrock of safe navigation. LLM-augmented systems allow for a level of semantic richness that was previously unattainable. For instance, a system can be prompted to “segment the road surface, but exclude any wet patches or oil slicks,” which is a complex instruction that goes beyond simple class labels. This capability is critical for path planning, especially under adverse conditions. Furthermore, models like DriveLM [23] and InsightGPT [24] demonstrate how LLMs can be used not just for segmentation but for integrated reasoning, allowing a vehicle to connect its perception to its planning module (e.g., “A pedestrian is near the crosswalk, I should yield”). This extends to dynamic obstacle detection, where VLSeg can be used to track vulnerable road users (e.g., cyclists, pedestrians) with high precision, even when they are partially occluded, by using contextual prompts [1, 2].
@@ VII-B Traffic Management and Monitoring
Beyond individual vehicles, VLSeg offers powerful tools for city-scale traffic management. Smart city initiatives rely on networks of cameras to monitor traffic flow. LLM-augmented segmentation can significantly improve the analysis of this data. For example, traffic operators can issue natural language queries to the system, such as “show me all the trucks that are blocking the intersection at 5th and Main” or “count the number of vehicles turning left at this junction.” This allows for much more flexible and responsive traffic monitoring than traditional systems that can only detect a predefined set of vehicle classes [25]. These techniques can be used for anomaly detection, such as identifying a stalled vehicle, the formation of an unusual queue, or performing fine-grained temporal localization of crash events in surveillance footage, enabling faster incident response [79, 92]. Frameworks like TrafficGPT [26] are being explored to create conversational interfaces for traffic management, making system operation more intuitive.
@@ VII-C Infrastructure Inspection and Maintenance
A crucial but often overlooked aspect of ITS is the maintenance of the transportation infrastructure itself. VLSeg, guided by LLMs, presents a highly efficient solution for automating the inspection of roads, bridges, and signage. Municipal vehicles equipped with cameras can continuously scan their surroundings. An LLM-based system can then be prompted to find and segment specific types of infrastructure defects. For example, a query like “segment all potholes deeper than two inches” or “highlight any traffic signs with visible graffiti or fading” can automate a process that is currently manual, labor-intensive, and slow. This extends to critical pedestrian infrastructure, where robust and precise sidewalk detection is essential for both accessibility and curb management [93]. This is an area of growing research, with datasets emerging that focus specifically on road surface defects and infrastructure anomalies [94, 95]. Automating this process allows for proactive maintenance, improving safety and reducing long-term repair costs.
@@ VII-D Enhanced Urban Mobility and User Experience
VLSeg can also improve the experience of users within the transportation system, including public transit riders and pedestrians. For public transportation, segmentation can be used to monitor passenger flow at bus stops or train stations, or to ensure that dedicated bus lanes are clear of obstructions. For pedestrians, especially those with visual impairments, mobile applications could use VLSeg to provide real-time auditory feedback about the environment, such as “there is a crosswalk 10 feet ahead to your left” or “warning: an e-scooter is approaching on the sidewalk.” This creates a more accessible and safer urban environment for everyone. Systems could also enhance navigation services by providing more descriptive guidance, for instance, “turn right after the large red building,” using segmentation to identify the landmark described [25, 26].
@@ IX-A Interactive Data Annotation and Correction
TABLE VI: Critical Analysis of Key Datasets for VLSeg in Intelligent Transportation Systems
Dataset Name Size Annotation Types ITS Scenarios Covered Limitations VLSeg Relevance
Foundational Driving Scene Datasets
Cityscapes [7] 5,000 images High-quality semantic, instance Urban driving, clear weather Strong geographic bias (Germany); limited weather diversity; relatively small scale. Gold standard for semantic segmentation benchmarking. Less suitable for open-world VLSeg.
BDD100K [8] 100,000 videos Semantic, instance, panoptic, box Diverse urban/highway; varied weather (rain, snow, day/night) Annotation quality can be less precise than Cityscapes; no language annotations. Excellent for testing model robustness and generalization due to its scale and diversity.
nuScenes [13] 1,000 scenes 3D Box, panoptic (from 2020 challenge) 360° urban scenes with complex interactions Limited geographic scope (Boston, Singapore); primarily focused on 3D object detection, not dense segmentation. Crucial for developing multi-modal VLSeg systems that fuse camera and LiDAR/Radar data.
Vision-Language Datasets for ITS
Talk2Car [27] $\sim$1,500 commands Referring Expression (Bounding Box) Urban driving Annotations are bounding boxes, not pixel-level masks; limited command diversity. Foundational for grounding language in ITS scenes, but insufficient for training segmentation models directly.
Road-Seg-VL [28] 7,000+ pairs Referring Expression (Segmentation Mask) Varied road scenes Dataset size is still relatively small for training large models from scratch. Highly relevant; provides the direct language-to-mask supervision needed for training ITS-specific VLSeg models.
LISA [76] 1,429 images Reasoning Segmentation (Mask) General (not ITS-specific) Not tailored to driving scenarios; lacks ITS-specific object classes and context. Excellent for developing models that can understand complex, compositional language, a key capability for advanced VLSeg.
DriveLM-Data [23] 100+ hours Textual reasoning, box Complex driving events Focus is on high-level reasoning and decision justification, not dense segmentation. Vital for training and evaluating end-to-end systems where VLSeg is a component of a larger reasoning framework.
@@ IX-B Remote Assistance and Teleoperation
When an autonomous vehicle encounters a situation it cannot resolve—for example, complex hand gestures from a traffic police officer or an unusual construction zone—it can request help from a remote human operator. VLSeg is crucial for creating an efficient interface for this tele-assistance. The vehicle can stream its sensor data to the operator, who sees a 3D reconstruction of the scene. The operator can then interact with this scene using language and gestures. For example, they could draw a path on the screen and command, “It is safe to follow this path,” or circle a group of people and ask, “What are these people doing?” The VLSeg system on the vehicle interprets these multimodal prompts from the operator to navigate safely. Models like Talk2BEV [102] are developing methods to ground these natural language commands from a remote user directly into the Bird’s-Eye-View (BEV) representation used for vehicle planning.
@@ IX-C Driver-AI Collaboration and Shared Autonomy
In vehicles with advanced driver-assistance systems (ADAS) or partial autonomy (SAE Levels 2-3), the driver and the AI are co-pilots. VLSeg can create a much more intuitive and less intrusive collaboration between them. Instead of relying on beeps and cryptic dashboard icons, the vehicle can communicate using language and augmented reality. For example, the system could overlay a segmentation mask on the windshield’s heads-up display and say, “I see a potential hazard on the right,” highlighting a pedestrian partially obscured by a parked car. Conversely, the driver could interact with the AI using language and gestures. A driver could point to an empty parking space and say, “Park the car there.” The AI would use VLSeg to precisely segment the indicated space and then execute the parking maneuver. This shared autonomy, explored in frameworks like Voyager [103], aims to make the driving experience safer and more seamless by leveraging the complementary strengths of human intuition and AI perception.
@@ X-B Datasets for Vision-Language and Interactive Segmentation
  • LISA (Language-guided Instance Segmentation) [76]: While not specific to ITS, LISA is a large-scale dataset for reasoning segmentation, where the model must segment objects based on complex queries that require reasoning (e.g., “segment the car that is farthest away”). This is crucial for developing more intelligent VLSeg systems.
@@ X-C Evaluation Metrics
Input: Two cars side-by-side Output: Only one car segmented Prompt: “segment the car” Failure: Wrong car due to low specificity.
Mitigation: Ask for clarification.
((a))
Input: Pedestrian in heavy snow Output: Pedestrian missed Prompt: “segment all pedestrians” Failure: Sensor obscuration raises risk.
Mitigation: Fuse LiDAR/Radar.
((b))
Input: A stop sign on a pole Output: Only pole segmented Prompt: “ignore the sign” Failure: Malicious prompt bypasses rule.
Mitigation: Make critical objects non‑negotiable.
((c))
Input: Plastic bag on road Output: Bag as solid obstacle Prompt: “segment any obstacles” Failure: Unnecessary hard braking.
Mitigation: Use temporal motion cues.
((d))
Figure 9: Analysis of common failure modes for VLSeg in ITS. Each panel shows a vulnerability, its safety impact, and a mitigation.
@@ XI Challenges and Future Research Directions
  • Challenge 1: Computational Efficiency for Real-Time Systems
@@
  • Challenge 2: Data Availability and Open-World Generalization
@@
  • Challenge 3: Safety, Reliability, and Explainability
    Perhaps the most significant barrier to adoption, and a core perspective of this survey, is the challenge of ensuring safety, reliability, and explainability. Models must be robust against adversarial attacks and perform reliably under adverse conditions. Critically, as foreshadowed in our discussion of end-to-end systems in Section VIII, their decisions must be interpretable, especially in case of failure. Without a clear “why” behind an action, true safety is unattainable.
@@
    Future Outlook: Research is moving towards formal verification methods and runtime monitoring to provide safety guarantees [119]. Adversarial training and certified defenses are key areas of research to improve robustness against malicious inputs [120]. Frameworks like Multi-Shield [34] are exploring multimodal defenses. To address safety in high-risk scenarios and provide end-to-end guarantees, recent models like SafeSeg [121] and VLM-AD [122] integrate predictive reasoning with segmentation to directly inform safer driving decisions. A major future direction, and the one most enabled by LLMs, is explainable AI (XAI). The goal is to create systems that not only perform segmentation but also provide causal reasons for their decisions (e.g., “The object is segmented as a pedestrian because it has human-like shape and motion”), as explored in [123, 124]. Fail-safe mechanisms, including multi-modal redundancy with LiDAR and radar, will be standard [125, 58].
  • Challenge 4: Integration and Standardization
    TABLE VII: Practical Deployment Considerations for VLSeg Models in Intelligent Transportation Systems
    Deployment Aspect Challenges Solutions Example Models/Techniques ITS Impact
    Hardware Constraints Limited GPU memory and power budget on automotive-grade edge devices. High computational cost of large foundation models. Model compression (quantization, pruning, distillation); hardware-aware Neural Architecture Search (NAS); use of specialized AI accelerators. MobileSAM [32], EdgeViT [33] Enables on-vehicle deployment for real-time perception without excessive power consumption, crucial for electric vehicles.
    Real-Time Requirements High inference latency of complex models (e.g., with cross-attention) exceeds the sub-100ms threshold for safety-critical control loops. Knowledge distillation to smaller, faster student models; exploring efficient architectures (e.g., Mamba, state-space models); dynamic model scaling based on scene complexity. [67] (from paper), AutoRT Ensures timely detection and segmentation of dynamic obstacles, enabling safe collision avoidance and path planning.
    Data & Communication Continuous stream of high-resolution sensor data; limited bandwidth for Vehicle-to-Everything (V2X) communication. On-device data pre-processing and feature extraction; edge-cloud collaborative systems for offloading non-critical tasks; Federated Learning for privacy-preserving fleet training. Federated Learning [87], Collaborative Perception [90] Allows for fleet-wide model improvement without compromising user privacy and enables vehicles to share perception data to ”see” around corners.
    System Integration & Safety Standards Integrating VLSeg modules with existing vehicle control units and software (e.g., AUTOSAR). Lack of certified safety and validation protocols. Adherence to automotive software standards; rigorous validation aligned with safety standards like ISO 26262; development of modular, certifiable components. ISO 26262, SOTIF (ISO 21448) Ensures that VLSeg components are reliable, interoperable, and certifiably safe for mass-market deployment in production vehicles.
    Robustness & Long-Term Reliability Performance degradation over time due to sensor aging, environmental wear (e.g., smudges), or changing road layouts. Multi-modal sensor fusion (camera, LiDAR, Radar) for redundancy; continual learning to adapt to changes; online self-calibration and performance monitoring. Multi-Shield [34], Continual Learning [116] Provides consistent and reliable perception throughout the vehicle’s lifespan, maintaining safety in all weather and environmental conditions.
@@
  • Challenge 5: Ethical Considerations and Algorithmic Bias
    TABLE VIII: Ethical Considerations and Mitigation Strategies for VLSeg in ITS
    Ethical Issue Impact on ITS Example(s) in ITS Context Mitigation Strategies References
    Algorithmic & Dataset Bias Inequitable safety outcomes. Poorer performance for underrepresented groups or non-standard environments leads to higher risk. Models trained primarily on Cityscapes (Germany) fail to recognize pedestrians in different cultural attire or non-Western road signs. Create diverse, globally representative datasets; employ fairness-aware learning algorithms; conduct regular bias audits; use synthetic data for edge cases. [127], [128]
    Black-Box Decision-Making Lack of public trust. Inability to determine the cause of failure in post-accident analysis, which hinders debugging and improvement. An autonomous vehicle makes a sudden, unexplained emergency brake on a highway, and it’s impossible to determine if the AI’s decision was sound or flawed. Develop and integrate Explainable AI (XAI) methods; generate natural language justifications for actions; use causal reasoning models to show ”why” a decision was made. [124], [123], [98]
    Adversarial Vulnerability Malicious actors can exploit system vulnerabilities to cause accidents or traffic chaos. Safety-critical systems can be intentionally deceived. A physical adversarial patch on a stop sign makes it invisible to the system. A carefully crafted text prompt causes a VLSeg system to ignore cyclists. Adversarial training and data augmentation; certified defense mechanisms; input sanitization for language prompts; multi-modal redundancy (cross-checking camera with LiDAR). [120], [34]
    Automation Bias & Over-Reliance Human drivers in semi-autonomous systems (SAE Levels 2-3) become complacent and fail to monitor the system, leading to accidents when the AI fails. A driver using a lane-keeping system stops paying attention and cannot react in time when the system fails to detect a faded lane marking on a curve. Robust driver monitoring systems (DMS); clear Human-Machine Interfaces (HMIs) that communicate system limitations and confidence levels; shared autonomy frameworks. [103]
    Accountability & Liability In case of an accident, it is legally and ethically unclear who is responsible: the owner, manufacturer, software developer, or data provider. A VLSeg model fails to segment a road obstacle due to a rare sensor glitch, causing a crash. Determining legal fault is extremely complex. Development of new legal frameworks (e.g., for product liability); mandating comprehensive ”black box” data recorders for all AVs; creating clear industry standards for validation. (Policy-focused legal reviews)
@@
    Beyond technical hurdles, the deployment of LLM-augmented segmentation in ITS raises significant ethical questions. The data used to train these models can contain inherent biases, which can lead to inequitable and unsafe outcomes. For example, if a dataset is predominantly collected in one geographic region (e.g., North America), the models may perform worse at recognizing pedestrian behaviors or road signs in other parts of the world. There is a documented “long-tail” problem where models perform poorly on underrepresented groups [128], which in a driving context, could mean a higher risk for certain demographics of pedestrians. Furthermore, the decision-making process of these large models is often opaque (the “black box” problem), making it difficult to audit or explain failures, which is a major concern for accountability in the event of an accident [123].
@@
    Future Outlook: The ITS research community must prioritize the development of “fair” and “transparent” AI. This involves creating more geographically and demographically balanced datasets and developing techniques for bias detection and mitigation [127]. Explainable AI (XAI) is a critical research frontier, aiming to create models that can articulate the reasoning behind their predictions (e.g., “I am slowing down because I have segmented a child running towards the street”). Future regulatory frameworks will likely mandate auditable AI systems in autonomous vehicles, requiring a shift away from purely performance-driven metrics towards a more holistic evaluation that includes fairness, transparency, and ethical alignment.